# Tracking the Feature Dynamics in LLM Training: A Mechanistic Study

### v2

@@ Abstract
Understanding training dynamics and feature evolution is crucial for the mechanistic interpretability of large language models (LLMs). Although sparse autoencoders (SAEs) have been used to identify features within LLMs, a clear picture of how these features evolve during training remains elusive. In this study, we: (1) introduce SAE-Track, a novel method to efficiently obtain a continual series of SAEs; (2) mechanistically investigate feature formation and develop a progress measure for it ; and (3) analyze and visualize feature drift during training. Our work provides new insights into the dynamics of features in LLMs, enhancing our understanding of training mechanisms and feature evolution.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Illustration of the SAE-Track Framework. Sparse Autoencoders (SAEs) are trained on residual stream activations from sequential training checkpoints (checkpoint[$k$]). Each SAE[$k$] is initialized using parameters from the previously trained SAE[$k-1$]. This approach ensures the continuity of the SAE across LLM checkpoints, Real-Time Training aligned with the availability of checkpoints, and Efficient Convergence, significantly reducing training steps.
@@
As LLMs increase in size and complexity, understanding their internal mechanisms has become a critical challenge. One notable phenomenon is polysemanticity, where individual neurons respond to mixtures of seemingly unrelated inputs, which complicates interpretability. This phenomenon is hypothesized to arise from the superposition of features, as demonstrated in (Elhage et al., 2022), which is based on the linear representation hypothesis (Park et al., 2023). A prominent approach to addressing polysemanticity in LLMs involves using SAEs (Bricken et al., 2023; Templeton et al., 2024). SAEs provide a dictionary learning framework to disentangle overlapping features in activation spaces, enabling the identification of more interpretable features. Studies have shown that SAEs can reveal interpretable representations (Bricken et al., 2023; Cunningham et al., 2023), scale to larger models such as Claude 3 Sonnet (Templeton et al., 2024), uncover structures in representations (Li et al., 2024), and analyze non-linear representations (Engels et al., 2024).
@@
Recent work has compared features across different settings—such as different layers (Jack Lindsey et al., 2024; Balagansky et al., 2024), between base and fine-tuned models (Jack Lindsey et al., 2024; Connor Kissane et al., 2024; Taras Kutsyk, 2024), and across different model scales (Lan et al., 2024). However, training dynamics remain underexplored, despite their importance in understanding the mechanisms underpinning LLM behavior.
There are some mechanistic interpretability studies related to training dynamics, but they mostly focus on specific abilities: (Olsson et al., 2022) linked the emergence of in-context learning ability to the development of induction heads; (Nanda et al., 2023) analyzed grokking and its relationship to generalization; (Qian et al., 2024) studied the trustworthiness development of LLMs during pre-training; and (Li et al., 2023) explored the formation of topic structures throughout training. However, using SAEs to study training dynamics offers a new perspective on the evolution of feature semantics and geometry throughout the entire training process.
In this paper, we address these gaps by conducting a comprehensive mechanistic analysis of feature evolution during the training of LLMs. Specifically, we propose SAE-Track and use it to study how features develop and stabilize in semantic meaning, and how they undergo directional changes. By doing so, we uncover novel insights into the evolution of feature semantics and geometry during training.
Our contributions are summarized as follows:
  • SAE-Track: We introduce a novel method for efficiently tracking feature evolution in LLMs through a continual SAE series across training checkpoints. SAE-Track enables detailed and reliable analysis of feature dynamics (Sec. 2).
@@
  • Feature Semantic Evolution Phases and Patterns: Using SAE-Track, we identify three distinct phases of feature semantic evolution—Initialization & Warmup, Emergent, and Convergent—as well as three primary transformation patterns: Maintaining, Shifting, and Grouping (Sec. 3).
  • Investigation of Feature Formation: We define the concept of feature regions and study feature formation based on this foundation. Furthermore, we introduce the Progress Measure to quantify the evolution of features (Sec. 4).
  • Analysis of Feature Drift: We analyze feature drift, focusing on the directional evolution and trajectories of features. For a specific feature, its direction is the normalized decoder vector, while its trajectory is the decoder vector across training checkpoints, both of which capture the directional evolution of features. Our study reveals that feature initially exhibit significant drift, continue to drift even after features are considered “formed,” and ultimately stabilize to their final state (Sec. 5).
  • Experiments on Models of Different Scales: We validate our findings using open-source LLM checkpoints, including Pythia (Biderman et al., 2023) and Stanford CRFM Mistral GPT-2 (Karamcheti et al., 2021), with models of varying scales (Pythia: 160M, 410M, 1.4B; Stanford CRFM GPT-2: Small, Medium), demonstrating the generality and scalability of our approach.
@@ 2 SAE-Track: Getting a Continual Series of SAEs
2.1 Preliminaries: Sparse Autoencoders
We follow the formulation in (Bricken et al., 2023; Templeton et al., 2024), where $\mathbf{x}\in\mathbb{R}^{D}$ denotes residual stream activations ($\mathbb{R}^{D}$ represents the activation space). This space is decomposed into $\mathbb{R}^{F}$(defined as feature space), where $F$ is the dictionary size of the sparse autoencoder.
The SAE reconstructs activations as:
where $\mathbf{W}^{\text{dec}}\in\mathbb{R}^{D\times F}$ are decoder weights, $\mathbf{b}^{\text{dec}}\in\mathbb{R}^{D}$ are biases, and $f_{i}(\cdot)$ represents feature activations computed as:
Here, $\mathbf{W}^{\text{enc}}\in\mathbb{R}^{F\times D}$ are encoder weights, $\mathbf{b}^{\text{enc}}\in\mathbb{R}^{F}$ are biases, and $c\in\{0,1\}$ is a constant determining whether the decoder bias is included in the input transformation.
The loss function $\mathcal{L}$ combines reconstruction error and L1 norm regularization:
where
$\mathcal{L}_{1}=\begin{cases}\sum_{i}f_{i}(\mathbf{x})\quad\text{(unit norm decoder constraint)},\\ \sum_{i}f_{i}(\mathbf{x})\cdot\|\mathbf{W}^{\text{dec}}_{:,i}\|_{2}\quad\text{(no decoder constraint)}.\end{cases}$ (4)
2.1.1 Key Intuitions
@@
Assume (1) the gradient norm $\|\nabla_{\Theta^{(<l)}}\mathcal{L}(\Theta)\|\leq G$ is bounded, and (2) the function $F^{(l)}$ is Lipschitz continuous with respect to $\Theta^{(<l)}$, with constant $L$. If the learning rate satisfies $\eta<\tfrac{\epsilon}{LG}$, then the activations evolve incrementally:
$\left\|\mathbf{x}^{(l,t)}-\mathbf{x}^{(l,t-1)}\right\|<\epsilon.$ (6)
@@
The proof is provided in Appendix A. Although real-world training may deviate from this idealized assumption, activation continuity motivates our approach to efficiently track feature evolution using a continual series of SAEs.
checkpoint 0 checkpoint 15–21 checkpoint 153 (final)
Maintaining
4/61632
1.  … relevant information?
2.  for the collection of status information
3.  the information generated
$\Rightarrow$ 1.  provide additional information about a response.
2.  providing information on walking, cycling
$\Rightarrow$ 1.  use information received from other agencies.
2.  granted access to personally identifiable information.
Grouping
4/36072
1.  Railway infrastructures
2.  civic coffers, infrastructural development, …
3.  infrastructural growth
4.  0090688-Kimberling1
5.  Kim and Eminem …
$\Rightarrow$ 1.  a compelling insight
2.  The Bottom Line
3.  … and its uneasy
4.  compelling page-turner
5.  N, KM, AK and AM are employees of …
$\Rightarrow$ 1.  address some of the most compelling questions …
2.  a credible source for …
3.  no credible reason to …
4.  a compelling proposition
5.  serious and credible
Shifting
[to concept-level]
4/30867
1.  Cataldo and De Marchi
2.  republic of Catalonia
3.  … paid to Cataldo,
4.  Serpico and Cataldo,
5.  Catalyst-Accelerated
6.  Graduate Catalogs
$\Rightarrow$ 1.  this Catalan government
2.  Frances Dunn. Louisa
3.  the Catalan parliament
4.  majority of Louisianans
5.  revolutionary leader Louis de Saint-Just…
$\Rightarrow$ 1.  Miguel Aleman is located
2.  remuneration from Luis
3.  … RTA), and Jose Fernando Covian – Nares
4.  writer Luis Pardo
5.  Fernando Torres scored
Shifting
[to token-level]
4/40671
1.  media is fueled by similar ignorance
2.  He claims ignorance
3.  either through ignorance or internet
4.  showing your ignorance
$\Rightarrow$ 1.  (astNode.isFunction)
2.  ta toute la fine graisse du dessus …
3.  analysis.PolynomialFunctionLagrangeFormTest)
4.  le pot de graisse se trouvait
$\Rightarrow$ 1.  @prop Function is Horizontal – returns
2.  commons.math.FunctionEvaluationExceptionTest)
3.  fn{Function |String}
[…]
Table 1: Feature Semantic Evolution Patterns. Illustration of the three primary feature evolution patterns during training: Maintaining, where features consistently activate for specific tokens across checkpoints; Grouping, where previously noisy features gradually organize into structured semantic ones; and Shifting, where a token-level feature either transitions to another token-level feature or evolves into a concept-level feature. Since concept-level features do not exist at initialization, Shifting exclusively begins from token-level features. The notation “layer/feature index” refers to the layer (preceding the residual stream) and the specific feature highlighted in the examples, with all tracking conducted on Pythia-410m-deduped.
Figure 2: Left: A toy example illustrating feature regions with different activation levels in a 2D activation space. The colored regions represent varying activation intensities for a single feature, emphasizing the distinction between activation levels. Right: Faithful feature regions for three features, visualizing how features partition the activation space into semantically meaningful regions. This highlights the role of SAE in isolating and identifying these regions.
Figure 3: UMAP Visualization. UMAP illustrates the evolution of activation sets for various features across Pythia-410m-deduped checkpoints, with each colored point representing a datapoint associated with a distinct feature. Circular markers indicate token-level features, while diamond-shaped markers represent concept-level features. The visualization highlights the transition from random activations at the start of training (Checkpoint 0) to semantically coherent clusters at later stages (Checkpoint 153) and showcases the distinct dynamics of the two feature types.
2.1.2 Methodology: Recurrent Initialization
To efficiently track feature evolution across training checkpoints, we propose SAE-Track, a framework that leverages recurrent initialization to construct a continual series of SAEs. Specifically:
  • Each SAE[$k$] is initialized with the parameters of the previously trained SAE[$k-1$], ensuring smooth transitions and preserving continuity in SAE parameter.
  • The process begins with $\text{SAE}[1]$, which is randomly initialized and trained on activations from the first checkpoint ($\text{checkpoint}[1]$). Subsequent SAEs are then trained sequentially using activations from their respective checkpoints.
As illustrated in Fig. 1, this method ensures that the decoder weights evolve incrementally across checkpoints, significantly reducing the training cost for subsequent SAEs—$\leq\tfrac{1}{20}$ of the initial training steps. Furthermore, it enables real-time training, where SAEs are trained as checkpoints become available, without requiring access to all checkpoints upfront. This means that SAE-Track can be trained in parallel with the training of LLMs, enabling real-time interpretation and supervision throughout the training process. It also maintains the characteristics of normally trained SAEs, ensuring that the learned representations are consistent with our commonly used approach (see Appendix F).
@@ 3 Semantic Evolution of Features
In this section, we examine the features in the SAE series generated by SAE-Track and analyze the patterns of semantic evolution exhibited by these features.
Features extracted via SAEs are categorized into two types based on their activation behavior. We will discuss their behavior in detail after providing the following definitions:
@@
A concept-level feature activates across a set of tokens that are semantically related to a broader concept. For instance, “authentication” and “getRole()” may activate for the concept “user authentication.”
Token-level features exist at all checkpoints, while concept-level features gradually emerge during the LLM’s learning process. Specifically, the evolution of features can be described in three stages:
  • Initialization and Warmup. Token-level features are present from the beginning, while other features remain as noise. At this stage, features lack meaningful associations beyond individual tokens, similar to the phenomenon described in (Bricken et al., 2023).
  • Emergent Phase. Concept-level features begin to emerge, representing abstract concepts, while token-level features persist in parallel. Meanwhile, noise features gradually vanish.
  • Convergent Phase. Both token-level and concept-level features converge into interpretable states, with concept-level features exhibiting coherent and semantically meaningful groupings.
We examine the semantic behavior of specific features throughout the LLM learning process. (Dead features and ultra-low activation features are excluded from this analysis, as well as all subsequent mechanistic analyses, as discussed in Appendix B.) Three distinct patterns of feature transformation are observed, as illustrated in Table 1:
  • Maintaining. Token-level features persist across checkpoints, maintaining consistent activations for the same token.
  • Shifting. Token-level features transition to different token-level ones or evolve into concept-level features, reflecting the reorganization of representations during training. This shifting process is inevitable for some features, as discussed in Appendix E.
  • Grouping. Previously noisy features gradually organize into structured semantic ones, forming distinct token- or concept-level features.
@@ 4 Analysis of Feature Formation
In this section, we analyze how features transition from noisy activations to meaningful representations throughout training. We formally define key concepts, track the evolution of feature regions, and propose a Progress Measure to quantitatively assess this process in both activation and feature spaces.
4.1 Understanding Feature Formation from a Geometric Perspective
Existing studies often emphasize the final state of training or assume that features are inherently monosemantic (Bricken et al., 2023; Templeton et al., 2024; Elhage et al., 2022). However, training a SAE at any model checkpoint reveals features that define separable regions in the activation space. While the specific properties of these features may vary across checkpoints, they can all be understood from a unified geometric perspective: they represent distinct regions within the activation space.

### v3

@@ Abstract
Understanding training dynamics and feature evolution is crucial for the mechanistic interpretability of large language models (LLMs). Although sparse autoencoders (SAEs) have been used to identify features within LLMs, a clear picture of how these features evolve during training remains elusive. In this study, we (1) introduce SAE-Track, a novel method for efficiently obtaining a continual series of SAEs, providing the foundation for a mechanistic study that covers (2) the semantic evolution of features, (3) the underlying processes of feature formation, and (4) the directional drift of feature vectors. Our work provides new insights into the dynamics of features in LLMs, enhancing our understanding of training mechanisms and feature evolution. For reproducibility, our code is available at https://github.com/Superposition09m/SAE-Track.
@@ 1 Introduction
As LLMs increase in size and complexity, understanding their internal mechanisms has become a critical challenge. Polysemanticity – where single neurons respond to mixtures of unrelated inputs – complicates interpretability. Sparse Autoencoders (SAEs) offer a promising approach to this challenge by disentangling overlapping features, enabling the extraction of more interpretable structures from activations bricken2023towards ; cunningham2023sparseautoencodershighlyinterpretable . Recent work demonstrates that SAEs can scale to models like Claude 3 Sonnet templeton2024scaling , uncover geometric structures in representations li2024geometryconceptssparseautoencoder ; engels2024not , and even intervene in model behavior chalnev2024improvingsteeringvectorstargeting ; yin2024direct ; farrell2024applyingsparseautoencodersunlearn .
@@
Researchers have also compared features across different settings, such as different layers crosscoder ; balagansky2024mechanisticpermutabilitymatchfeatures , between base and fine-tuned models crosscoder ; saeusuastransfer ; saesbasefinetuned , and across different model scales lan2024sparseautoencodersrevealuniversal . However, LLMs’ training dynamics remain underexplored, despite their importance in understanding the mechanisms underpinning LLM behavior.
While prior work in mechanistic interpretability has shed light on individual phenomena related to training dynamics – for instance, olsson2022context on induction head formation and in-context learning, nanda2023progress on grokking phenomenon of generalization, qian2024towards on trustworthiness emergence in LLM pre-training, NEURIPS2024_5aa96d1c on interactions, and li2023transformers on topic-structure development – these studies often remain narrowly scoped, focusing on isolated abilities or providing static snapshots of network behavior. What is largely missing, therefore, is a unified and continuous framework that systematically monitors how fundamental feature representations evolve – both semantically and geometrically – throughout the entire training trajectory.
To address this critical gap, this paper introduces SAE-Track to effectively track feature evolution and conduct a mechanistic study based on this framework, covering the semantic evolution of features, feature formation, and feature vector drift analysis.
Our main contributions are as follows:
  • SAE-Track: A novel method for efficiently obtaining a continual series of SAEs, providing the foundation for a detailed mechanistic study of feature evolution in LLMs (Sec. 2).
@@
  • Semantic Evolution Phases and Patterns: Identification of three characteristic phases – Initialization & Warmup, Emergent, and Convergent – and three primary transformation patterns: Maintaining, Shifting, and Grouping, capturing the gradual transition from a randomized transformer to a well-trained one (Sec. 3).
  • Feature Formation: Geometric modeling of feature formation as the convergence of datapoints into localized regions, with a novel Progress Measure that captures both the gradual nature of this process and the distinct dynamics of token-level and concept-level features (Sec. 4).
  • Feature Drift and Trajectory Analysis: Analysis of feature drift, revealing that feature directions undergo significant, three-phase adjustments. Counter-intuitively, this drift persists even after features are semantically “formed”, with full stabilization only occurring late in training (Sec. 5).
  • Extensibility Validation: Extensive experiments with open-source LLMs (Pythia and Stanford CRFM GPT-2) of varying scales (124M to 1.4B) and across different residual stream layers, confirming the extensibility and generality of our approach. (Pythia-410M-deduped, layer 4 for main paper results; additional models and layers in Appendix K.) We have tested our method on as many feasible models as possible within our computational constraints and available checkpoints (See limitations discussed in Appendix A.)
@@ 2 SAE-Track: Getting a Continual Series of SAEs
2.1 Preliminaries: Sparse Autoencoders (SAEs)
We employ Sparse Autoencoders (SAEs) bricken2023towards ; templeton2024scaling , an unsupervised method to learn sparse, interpretable features from data such as LLM activations $\mathbf{x}\in\mathbb{R}^{D}$. An SAE is trained to reconstruct $\mathbf{x}$ as $\widehat{\mathbf{x}}$ from a parsimonious set of learned dictionary features weighted by their activations $f_{i}(\mathbf{x})$. This is achieved by minimizing the loss function $\mathcal{L}$ which combines reconstruction error with an L1 sparsity penalty on $f_{i}(\mathbf{x})$:
$\mathcal{L}=\mathbb{E}_{\mathbf{x}}\left[\|\mathbf{x}-\widehat{\mathbf{x}}\|_{2}^{2}+\lambda\mathcal{L}_{1}\right].$ (1)
The detailed mathematical formulation of the SAE architecture, including the computation of feature activations $f_{i}(\mathbf{x})$, the reconstruction $\widehat{\mathbf{x}}$, and specific forms of the L1 penalty $\mathcal{L}_{1}$, is provided in Appendix B.
2.2 Key Intuitions
@@
Figure 1: SAE-Track Framework. Sparse Autoencoders (SAEs) are trained on residual stream activations from sequential LLM checkpoints. Each SAE[k] is initialized from the previous SAE[k-1], enabling continual tracking, real-time parallel training, and efficient computation with reduced steps (e.g., <1/20) for subsequent SAEs.
@@
Assume (1) the gradient norm $\|\nabla_{\Theta^{(<l)}}\mathcal{L}(\Theta)\|\leq G$ is bounded, and (2) the function $F^{(l)}$ is Lipschitz continuous with respect to $\Theta^{(<l)}$, with constant $L$. Let $\epsilon$ be a small positive constant. If the learning rate satisfies $\eta<\tfrac{\epsilon}{LG}$, then the activations evolve incrementally:
$\left\|\mathbf{x}^{(l,t)}-\mathbf{x}^{(l,t-1)}\right\|<\epsilon,$ (3)
@@
The proof is provided in Appendix C. This theorem formalizes the intuition that, under bounded gradients and a suitably small learning rate, each parameter update induces only a minor shift in layer activations. Although real-world training may deviate from this idealized assumption, activation continuity motivates our approach to efficiently track feature evolution using a continual series of SAEs.
2.3 Methodology: Recurrent Initialization
To efficiently track feature evolution, we propose SAE-Track, which constructs a continual series of SAEs via recurrent initialization: each $\text{SAE}[k]$ is initialized from $\text{SAE}[k-1]$ (with $\text{SAE}[1]$ randomly initialized) and trained on activations from $\text{checkpoint}[k]$. As illustrated in Fig. 1, SAE-Track is designed to be continual, real-time, and efficient. It maintains feature continuity by leveraging recurrent initialization, supports real-time training alongside LLM checkpoints, and significantly reduces training cost by requiring fewer training steps for subsequent SAEs.
To validate the effectiveness of this approach, we conducted a series of comparative experiments (see Appendix I.1 and Appendix I.2). Appendix I.1 demonstrates that SAE-Track not only significantly accelerates convergence but also produces individual SAEs that exhibit similar behavior to those trained using conventional methods, ensuring that the learned representations remain consistent with standard training approaches. Appendix I.2 further confirms that the observed feature convergence is not merely an artifact of the SAE-Track process, but a fundamental property of the model’s feature dynamics.
Figure 2: Feature transition patterns.(on Pythia-410m.) The left panel displays examples of feature transitions—maintaining, grouping (into concept/token-level feature), and shifting (to concept/token-level feature)—across training checkpoints. The right panel shows the statistics of each transition type (on 256 random sampled features).
@@ 3 Semantic Evolution of Features
This section delves into the semantic evolution of features as LLMs undergo training, leveraging the series of SAEs generated by SAE-Track. Our analysis method involves tracking the interpretation of fixed feature indices (IDs) across these sequential SAEs, which lead to our primary conclusion regarding the phases and patterns of semantic evolution:
Conclusion 1 Based on the taxonomy distinguishing between token-level and concept-level features—itself an observation from their differing activation behaviors (see Definition 3.1 and Definition 3.2)—we identify a characteristic three-phase process governing their semantic evolution throughout training. Furthermore, this distinction allows us to categorize the observed feature transformation patterns, as exemplified in Fig. 2, into three primary patterns.
The distinction between token-level and concept-level features is foundational to understanding their evolution. As evidenced by their differing activation patterns over the training examples shown in Fig. 2 (Left), we define them as follows:
@@
A concept-level feature activates across a set of tokens that are semantically related to a broader concept. For instance, “authentication” and “getRole()” may activate for the concept “user authentication.” This category also encompasses weak concept-level features, such as those based on morphological variants(“arrive” and “arrives”), as detailed in Appendix F.
Token-level features are typically present from initial checkpoints, while concept-level features emerge more gradually. This evolution, observable by comparing early versus late checkpoint examples in Fig. 2 (Left), can be described in three stages:
  • Initialization and Warmup. Token-level features are present from the outset, while other activations often appear as noise with limited semantic association beyond individual tokens, as seen in early checkpoint examples (e.g., “ckpt 0” in Fig. 2).
  • Emergent Phase. Concept-level features begin to form, with activations grouping around semantically related tokens, while token-level features still exist and noise features diminishes, visible in intermediate checkpoints (e.g., “ckpt 15-21” in Fig. 2).
  • Convergent Phase. Both feature types stabilize into interpretable states, with concept-level features forming coherent semantic groups by later checkpoints (e.g., “ckpt 153(final)” in Fig. 2).
Further examination of individual feature ID evolutions, by analyzing changes in their top-k activating examples as illustrated in Fig. 2 (Left), reveals three distinct transformation patterns. Their statistics, from a sample of 256 features, is shown in Fig. 2 (Right). These patterns are:
  • Maintaining. Token-level features often persist across checkpoints with consistent token activations, as exemplified by the first row of examples in Fig. 2 (Left).
  • Shifting. Some features alter their primary semantic association during training. These shifts can be towards a new token-level representation (exemplified in the fifth row of Fig. 2 (Left)) or an evolution into a broader concept-level one (exemplified in the fourth row of Fig. 2 (Left)), reflecting representational reorganization (see Appendix H).
  • Grouping. Features initially exhibiting noisy or diffuse activations gradually organize into semantically coherent structures. This can result in the formation of new concept-level features (exemplified in the second row of Fig. 2 (Left)) or new token-level features (exemplified in the third row of Fig. 2 (Left)).
Our proposed transformation patterns form a comprehensive taxonomy, since they encompass all observed semantic developmental pathways from initial (token-level or noise) to mature (token-level or concept-level) feature states.
While this qualitative characterization offers valuable insights, its interpretative nature (common in SAE analysis though) has inherent limitations. To go beyond these limitations and provide more robust validation, we conduct further quantitative mechanistic investigations in subsequent sections. Both Section 4 (on feature formation) and Section 5 (on feature drift) quantitatively support the three-phase development model, with the latter also providing a geometric complement to the semantic analysis herein.
@@ 4 Analysis of Feature Formation
This section analyzes feature formation, from noisy activations to meaningful representations. We first frame this geometrically as datapoint convergence into local regions. We then use a novel Progress Measure to quantitatively show this process is gradual, and to explain the distinct emergence and learning dynamics of token-level versus concept-level features.
Conclusion 2: Feature Formation Mechanisms Feature formation involves geometric convergence of activations into local regions. Our Progress Measure confirms this is gradual, elucidating token-level feature initial existence versus later concept-level feature learning, and supports a three-phase development model.
4.1 Geometric Perspective and Qualitative Illustration of Feature Formation
Existing studies often emphasize the final state of training or assume that features are inherently monosemantic bricken2023towards ; templeton2024scaling ; elhage2022toy . However, training a SAE at any model checkpoint reveals features that define separable regions in the activation space. While the specific properties of these features may vary across checkpoints, they can all be understood from a unified geometric perspective: they represent distinct regions within the activation space.